# Evolutionary Optimization of Physics-Informed Neural Networks: Evo-PINN Frontiers and Opportunities

### v3

@@ IV-D3 Meta-Learning Neural Architecture, Loss Function, and Optimizer
Most gradient-based meta-learning PINN approaches proposed in the literature aim to find an optimal initialization of the network weights. In the study by [192], a meta-learning framework for PINN loss function that incorporates the underlying physics of a parametrized PDE has demonstrated significant performance improvement for new tasks. Besides, a meta-learning gradient-based optimizer [193] has been recently introduced for improving hand-designed optimization rules. The meta-learned optimizer has shown significant improvement in physics-informed learning tasks compared to traditional fixed, hand-crafted optimizers, such as the Adam optimizer.
@@
There are already a handful of studies that have proposed using EAs to find performant neural architectures and hyperparameters for a single PINN; ref. to Subsection IV-B. There is huge scope for future research and development into using EAs to jointly meta-learn model initialization, neural architecture, loss function (hyperparameters), and even the gradient-based optimizer to drastically improve convergence on a set of PINN tasks. For example, a recent work proposes the use of an EA to jointly optimize the neural architecture and network weights for a convolutional neural network-type PINN. The meta-learned PINN model integrates a physics-informed pseudoinverse computation on the final CNN layer to ensure physics compliance during prediction of a new scenario. The proposed method demonstrated enhanced generalization across diverse tasks with varying initial conditions, PDE parameters, and geometries, significantly outperforming other data-driven models like U-Nets and Fourier Neural Operators.
@@
There exists a multitude of evolutionary computation techniques for achieving meta-learnable optimizers, loss functions, neural architectures, and model initializations over a distribution of physics-informed learning tasks. They can provide diverse training strategies for exploration to achieve effective and efficient training of a generalizable PINN. Lastly, we would like to make additional remarks on the important role of labeled data in meta-learning PINN. In general, the meta-learning of PINNs incurs a higher upfront training cost to achieve good generalization. Some proposed methods such as the Baldwinian-PINN cleverly utilize a data-centric meta-learning objective, i.e., using data MSE rather than purely physics-driven loss. Meta-learning PINN algorithms can be meaningfully guided by a few tasks with labeled data, easing optimization difficulties caused by the complex loss landscape defined by the physics-informed loss.
IV-D4 Other Approaches for Advancing Generalizability
A notable alternative to generalized PINNs is through neural operator networks [194, 195, 196], which, by learning mappings between infinite-dimensional function spaces, allow models to solve PDEs in real-time under many different ICs or BCs, requiring only light task-specific fine-tuning. However, unlike parameterization-agnostic Baldwinian-PINNs, neural operators call for explicit parameterization of the conditions over which the model is to generalize. Although this does make neural operators less flexible compared to meta-learning, they are considered an important complementary approach to model generalization. At the time of writing of this paper, the effectiveness of EAs for optimizing physics-informed neural operator nets remains an open question. Some foundational work bringing randomization into shallow operator networks has appeared in [197], showing their potential to outperform standard operator nets by several orders of magnitude in terms of approximation accuracy and computational cost.
@@ IV-E Evolving Interpretable Models Beyond PINNs
In contrast to neural nets, the subfield of genetic programming in evolutionary computation has a long history of successfully discovering free-form natural laws from data through symbolic regression [199, 200]. Recently, these algorithms have even been used for solving high-dimensional PDEs, with the distinct advantage of offering compact analytical expressions approximating complex solutions [201, 202, 203]. Balancing compactness and model accuracy is indeed a non-trivial task, one that is well-suited to evolutionary multi-objective optimization. In addition to symbolic regression, a highly interpretable mathematical description of dynamical systems can be obtained by identifying a best-fit linear dynamics model (i.e., a linear system of governing ODEs) from data. Although finding such a data-driven model that simultaneously satisfies physics constraints translates into a highly constrained, nonlinear and non-convex optimization problem [204], it has been shown to be efficiently solvable by evolution strategies [205]. Adaptations of this evolutionary approach to identify nonlinear dynamics, through neural ODEs [206], have also been explored [207], with the final prediction accuracy and training stability of evolution exceeding those of models trained by gradient descent.
@@ V Conclusion
However, like any new technique, PINNs inevitably call for new technical innovations. Many of the approaches commonly used in training standard data-driven neural nets need to be re-evaluated, such as the method of network initialization and the choice of activation functions. The use of purely gradient-based optimization, which includes the hugely successful family of SGD algorithms, is especially worthy of reconsideration in the light of analyses summarized in this work. Identified challenges pertaining to local minima traps, fixed point traps, and vanishing gradients of physics-informed learning objectives highlight how these can prove deleterious for PINNs. These obstacles to conventional gradient-based optimization serve to outline ways in which gradient-free algorithms of computational intelligence may be uniquely advantageous as substitutes to or in synergy with SGD. Indeed, the crafting of synergistic optimization frameworks that meld gradient-based and gradient-free optimization unveils novel research questions and directions, firmly positioning scientific machine learning as an emerging area of translational interest for the computational intelligence community.

### v5

@@ IV-D3 Meta-Learning Neural Architecture, Loss Function, and Optimizer
Most gradient-based meta-learning PINN approaches proposed in the literature aim to find an optimal initialization of the network weights. In the study by [187], a meta-learning framework for the PINN loss function that incorporates the underlying physics of a parametrized PDE has demonstrated significant performance improvement for new tasks. In [188], the complexity of the PINN optimization problem was proposed to be tackled by learning a customized optimizer that’s itself parameterized by a neural network. Notably, the learning of the optimizer was done using a persistent ES algorithm [189], and showed that meta-learned optimizers could substantially improve vanilla PINNs over standard hand-crafted optimizers such as Adam.
@@
There are already a handful of studies that have proposed using EAs to find performant neural architectures and hyperparameters for a single PINN (ref. to Subsection IV-B). There is huge scope for future research and development into using EAs to jointly meta-learn model initialization, neural architecture, loss function (hyperparameters), and even the gradient-based optimizer to drastically improve convergence on a set of PINN tasks. For example, an EA can jointly optimize the neural architecture and network weights for a PINN such that the meta-learned PINN model utilizes physics compliance during prediction of a new scenario to enhance generalization across diverse tasks with varying initial conditions, PDE parameters, and geometries.
@@
There is significant potential in exploring evolutionary computation methods to develop meta-learnable optimizers, loss functions, neural architectures, and initializations tailored to diverse physics-informed learning tasks, thereby presenting opportunities to enhance the generalizability of PINN models. Lastly, we would like to make additional remarks on the important role of labeled data in meta-learning PINN. In general, the meta-learning incurs a higher upfront training cost to achieve good generalization. Some proposed methods such as the Baldwinian-PINN cleverly utilize a data-centric meta-learning objective, i.e., using data MSE rather than purely physics-driven loss. Meta-learning PINN algorithms can be meaningfully guided by a few tasks with labeled data, easing optimization difficulties caused by the complex loss landscape defined by the physics-informed loss.
IV-D4 Neural Operator Networks
A notable alternative to learning general PDE solutions is the use of neural operator networks in a purely data-driven setting [190, 191, 192]. By learning mappings between infinite-dimensional function spaces, neural operators enable prediction of solutions in real-time under different PDE source terms, ICs or BCs. For instance, consider the 1D Poisson equation $\frac{d^{2}u}{dx^{2}}=f(x)$; a neural operator aims to learn the mapping $G:f\to u$ from the source term $f(x)$ to the PDE solution $u(x)$. Prominent neural operator architectures encompass the Deep Operator Network (DeepONet) [193] and the Fourier Neural Operator (FNO) [194]. DeepONet comprises a branch network to encode the input function at predetermined sample locations, and a trunk network to encode spatial coordinates at which the output function is evaluated. Conversely, FNO employs spectral convolutions to efficiently model nonlocal interactions and directly predict the entire field solution, making it particularly suited for structured-grid data [195]. Further extensions, such as graph-based neural operators [196], transformer operators [197], and geometry-informed operators [198], broaden this framework to manage irregular domains and bolster computational efficiency. Similar to other data-driven methods, neural operators are inherently data-intensive, typically necessitating large PDE simulation datasets to achieve robust generalization.
The integration of physics-informed losses, such as PDE constraints, into neural operators can substantially reduce training data requirements by leveraging known physics [199]. This approach establishes PINNs and neural operators as complementary techniques. Nonetheless, analogous to parameterized PINNs, adherence to physical laws during training does not necessarily translate to their preservation in out-of-sample predictions, thereby presenting common generalization challenges. As a result, neural operators often resort to using larger training datasets to enhance generalization. Furthermore, unlike parameterization-agnostic Baldwinian-PINNs, they demand explicit parameterization of the conditions under which the model is intended to generalize. Although this explicit requirement diminishes the flexibility of neural operators relative to meta-learning strategies, they remain a vital complementary method for generalized neural PDE solvers. Currently, the effectiveness of EAs for optimizing neural operators remains an open area of research. Early work bringing randomization into shallow operator networks has appeared in [200], showing their potential to outperform standard operator networks in both approximation accuracy and computational efficiency.
@@ IV-E Evolving Interpretable Models Beyond PINNs
In contrast to neural networks, the subfield of genetic programming in evolutionary computation has a long history of successfully discovering free-form natural laws from data through symbolic regression [202, 203]. Recently, these algorithms have even been used for solving high-dimensional PDEs, with the distinct advantage of offering compact analytical expressions approximating complex solutions [204, 205, 206]. Balancing compactness and model accuracy is a non-trivial task that is well-suited to evolutionary multi-objective optimization. In addition to symbolic regression, a highly interpretable mathematical description of dynamical systems can be obtained by identifying a best-fit linear dynamics model (i.e., a linear system of governing ODEs) from data. Although finding such a data-driven model that simultaneously satisfies physics constraints translates into a highly constrained, nonlinear and non-convex optimization problem [207], it has been shown to be efficiently solvable by evolution strategies [208]. Adaptations of this evolutionary approach to identify nonlinear dynamics have also been explored [209], with the final prediction accuracy and training stability of evolution exceeding those of models trained by gradient descent.
@@ V Conclusion
However, like any new technique, PINNs inevitably call for new technical innovations. Many of the approaches commonly used in training standard data-driven neural networks need to be re-evaluated, such as the method of network initialization and the choice of activation functions. The use of purely gradient-based optimization, which includes the hugely successful family of SGD algorithms, is especially worthy of reconsideration in the light of analyses summarized in this work. Identified challenges pertaining to local minima traps, fixed point traps, and vanishing gradients of physics-informed learning objectives highlight how these can prove deleterious for PINNs. These obstacles to conventional gradient-based optimization serve to outline ways in which gradient-free algorithms of computational intelligence may be uniquely advantageous as substitutes for or in synergy with SGD. Indeed, the crafting of synergistic optimization frameworks that meld gradient-based and gradient-free optimization unveils novel research questions and directions, firmly positioning scientific machine learning as an emerging area of translational interest for the computational intelligence community.
Acknowledgment
This research was in part supported by the National Research Foundation, Singapore through the AI Singapore Programme, under the project “AI-based urban cooling technology development” (Award No. AISG3-TC-2024-014-SGKR), in part by the National Research Foundation, Singapore and DSO National Laboratories under the AI Singapore Programme - “Design Beyond What You Know: Material Informed Differential Generative AI (MIDGAI) for LightWeight High-Entropy Alloys and Multi-functional Composites (Stage 1b)” (Award No. AISG2-GC-2023-010). Abhishek Gupta is supported by the Ramanujan Fellowship from the Anusandhan National Research Foundation, Government of India (Grant No. RJF/2022/000115).